# Mechanistic Modeling of Continuous Lyophilization for Pharmaceutical Manufacturing

### v1

@@ 3.2.5 Cooling
Thermal radiation is significantly more complicated in terms of modeling. The theories and model equations described below are mainly based on Refs. [31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization. Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; please refer to the aforementioned references for more details. Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively. In general, the net radiant energy receiving by surface 1 can be written as
$Q_{\textrm{rad}}=\sigma A_{1}\mathcal{F}(T_{2}^{4}-T_{1}^{4}),$ (45)
@@
where $A$ is the surface area, $\varepsilon$ is the emissivity, $\mathcal{F}$ is the transfer factor, and the subscripts ‘1’ and ‘2’ denote surfaces 1 and 2, respectively. The transfer factor is dependent on geometry and material type (emissivity). For any two diffuse, gray surfaces that form an enclosure,
$Q_{\textrm{rad}}=\dfrac{\sigma(T_{2}^{4}-T_{1}^{4})}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (46)
@@
where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 ($F_{1-2}=1$) and surface 2 is much larger than surface 1, (46) can be simplified as
$Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}(T^{4}_{2}-T^{4}_{1}).$ (47)
@@
In (47), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Now consider the case when surface 1 is the vial and surface 2 is the chamber walls. In this case, (47) becomes
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}(T^{4}_{\textrm{c}}-T^{4})$ (48)
for the side surface and
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{z}(T^{4}_{\textrm{c}}-T^{4})$ (49)
@@
for the top and bottom surfaces. Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$. The amount of radiative heat for the single-vial case represented by (48) and (49) defines the upper bound on the radiative heat that the vial can receive from the chamber walls. This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among vials. As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$. Therefore, (48) needs to be modified to account for interactions between vials, whereas (49) can be used directly because thermal radiation at the top and bottom surfaces is not influenced by the other vials.
@@
Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials. In batch lyophilization, vials are arranged as an array consisting of many rows, which results in significant differences in heat transfer conditions between the outer and inner vials. In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls. A rigorous way of modeling this complicated radiation is known as the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref. [39]. In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by Ref. [6]. As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials. Therefore, instead of using the radiation network approach, we modify (48) to follow (45) as
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}(T^{4}_{\textrm{c}}-T^{4}),$ (50)
@@
where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph. This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (46). This technique is similar to the hybrid approach presented in Ref. [39]. Note that this approximation is for the suspended-vial configuration, in which heat transfer between vials is not significant; otherwise, the radiation network approach is recommended.
@@
When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; for the temperature of about 300K, the error is about 0.1% [31]. The linearized equation can then be combined with the convection part, with the overall heat transfer coefficient taking into account of both convection and radiation, e.g., (34).
@@
The proposed model developed in Sections 3.2, 3.3, and 3.4 require a large number of input parameters. This section defines the default parameter values for our model to be used in the subsequent sections in this work. These parameter values are either obtained from the literature or set based on the typical values used in lyophilization. This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively. Table LABEL:Tab:DefaultParameters shows the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
@@
$\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix A)
@@
$T_{\textrm{c}}$ 273 (before VISF) K
@@
$R_{\textrm{p0}}$ 1.5$\times$104 m/s estimated from data in Ref. [6]
$R_{\textrm{p1}}$ 3.4$\times$107 1/s estimated from data in Ref. [6]
$R_{\textrm{p2}}$ 1 1/m estimated from data in Ref. [6]
@@
$H$ 7.2$\times$10-3 m calculated (see Appendix A)
$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix A)
$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix A)
@@
Parameters that should be estimated from experimental data include heat transfer coefficients, mass transfer coefficients and cake resistance, and desorption-related parameters (frequency factor and activation energy).
@@
In this article, the mechanistic model was implemented in MATLAB R2023a, and so all results and plots were generated using MATLAB. A system of ordinary differential equations (ODEs) was integrated with ode15s. Partial differential equations (PDEs) were discretized spatially to produce a system of ODEs, which was then fed to the ODE solver (see Appendix B for the detailed numerical methods).
@@
The mechanistic model presented in Section 3 is validated using the experimental data reported in Ref. [6]. The model is validated for each lyophilization step separately, with the parameters reported in Table 2. The default values defined in Table LABEL:Tab:DefaultParameters are used for parameters not shown in Table 2.
@@
$T_{\textrm{c}}$ 272 K estimated from data in Ref. [6]
@@
$T_{\textrm{b}}$ 275 K estimated from data in Ref. [6]
@@
$T_{\textrm{c}}$ 285 K estimated from data in Ref. [6]
@@
Figure 4: Model validation for the lyophilization of suspended vials using the experimental data from Ref. [6]. Panel A shows the model prediction and experimental data for the product temperature during the freezing step. Panel B shows the model prediction and experimental data for the product temperature (assumed to be measured at the bottom surface) during the primary drying step. The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively. Panel C shows the model prediction and experimental data for the bound water concentration during the secondary drying step. The initial concentrations for Cases 3a and 3b are 0.0864 and 0.0739 kg water/kg solid, respectively.
@@
For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig. 4A). During the preconditioning phase, the temperature decreases from its initial value of about 280 K to the nucleation temperature of about 263 K, where nucleation starts. At the first nucleation, the temperature instantaneously increases from 263 K to about the equilibrium freezing point of about 272 K. The temperature slowly decreases during the solidification phase, which takes about 30 min, following the freezing-point depression. Subsequently, the product temperature decreases and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
@@
The model can accurately predict the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig. 4B). The product temperature increases fast at the beginning because the driving force for sublimation is low, and so most of the heat input is used to increase the temperature. As time progresses, the driving force for sublimation is higher, resulting in most of the heat used for sublimation instead of increasing the product temperature. The maximum error in temperature prediction is about 2–3 K, while that of drying time prediction is about 0.2 for Case 2a and 0.4 for Case 2b ($\sim$ 6%). This error could be partly because of the uncertainty of cake resistance, which is caused by a stochastic freezing process.
@@
In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.0864 and 0.0739 kg water/kg solid). The concentration decreases exponentially following the linear driving force model, where the simulations are terminated when the concentration reaches 0.01 kg water/kg solid. It is important to note that the data used in Fig. 4C are based on the shelf temperature of 293 K. To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures such that the effects of temperature on the desorption process can be quantified.
Case Studies
The previous sections detail the development and validation of our mechanistic model for lyophilization. This section demonstrates various applications of our model via four different case studies.
Process design and optimization
@@
The first case study focuses on model-based design and optimization of the entire continuous lyophilization process. One of the critical aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production. In the context of continuous lyophilization, our model can be used to study the evolution of critical process parameters, e.g., temperature, moisture content, throughout the process at various operating conditions, which can help guide the process design and optimization.
@@
This work not only presents the first mechanistic model for continuous lyophilization, but also is one of the very few studies that proposes a complete model that includes all three steps in lyophilization, including freezing, primary drying, and secondary drying. This is important specifically for continuous manufacturing where the entire process should be designed and optimized simultaneously.
@@
By using the default parameters in Table LABEL:Tab:DefaultParameters, we simulate a complete continuous lyophilization cycle including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig. 5). The evolution of the product temperature (Fig. 5A), mass of ice (Fig. 5B), and bound water concentration (Fig. 5C) can be obtained, given the operating pressure (Fig. 5D) and temperature (Fig. 5E). This information could be useful for various purposes. For example, it can guide the process and equipment design concerning chamber sizing and vial moving speed to accommodate the residence time of each step. Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time. It is important to note that the entire simulation shown in Fig. 5 can be computed in less than 1 second on a normal laptop, which makes the model very practical to be implemented for any purposes.
@@
Figure 5: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) concentration of bound water, (D) operating pressure, and (E) operating temperature.
Studying vacuum-induced surface freezing
@@
Our model can be simulated for the whole lyophilization process as shown in Section 5.1 or simulated separately for each step in the lyophilization cycle to provide insights into some part of the process. This section employs the model to study the freezing step, which usually receives little attention compared to the drying steps.
The proposed model can simulate different freezing protocols, namely uncontrolled (aka spontaneous) and controlled nucleation. Spontaneous nucleation was already investigate in detail by Refs. [11, 12], and so here we consider the controlled nucleation case via VISF, which is preferable for continuous lyophilization. As an example, the evolution of the product temperature and amount of water evaporating during VISF at different pressure profiles is simulated (Fig. 6). Varying among different systems, this information could be used to design a freezing protocol, e.g., pressure profile, that achieves the desired nucleation temperature.
@@
Figure 6: Simulation results showing the evolution of the product during VISF at different pressure profiles. Panel A shows the pressure profiles that decrease linearly from $10^{5}$ Pa to $10^{4}$ Pa, 1,000, and 100 Pa, respectively. Panel B shows the evolution of the product temperature. Panel C shows the total mass of water, which implies the evaporation rate. External heat transfer with the wall and cryogenic gas is excluded so that only the contribution from VISF is considered.
Model modifications for conventional drying
@@
Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization. The major modifications are to turn off the thermal radiation part in the model and recalculate/estimate the heat transfer coefficient and cake resistance. In this case study, we consider batch primary drying and secondary drying using the data from Refs. [18] and [21], with the parameters reported in Table 3.
@@
$\mathcal{F}$ 0 turn off the radiation part
@@
$\mathcal{F}$ 0 turn off the radiation part
@@
    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 (Case 4b).
@@
    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264.09 K to 312 K for Case 5.
Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig. 7). The model can also effectively predict the product temperature and bound water concentration during the secondary drying step (Fig. 8). With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
Figure 7: Comparison between our model prediction and the experimental/simulation data from Ref. [18] for the primary drying step in conventional lyophilization. Panel A shows the model prediction and the experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and reference simulation data for the sublimation front position. The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15, while other conditions are identical.
Figure 8: Comparison between our model prediction and the experimental data from Ref. [21] for the secondary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and experimental data for the concentration of bound water.
@@
Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily at its desired conditions. A better process model should be able to simulate important abnormal conditions, e.g., equipment failure, that could occur due to various unexpected scenarios, which is critical in continuous manufacturing as these abnormal operations could affect the reliability, availability, and maintainability (RAM) as well as the safety of the plant/process.
In lyophilization, one of the crucial design considerations is associated with a condenser. During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1]. This is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase. This scenario may not lead to safety-related issues because the operating pressure in lyophilization is usually low and the amount of ice/water is not high enough to create overpressure. Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (31)), which then prolongs the primary drying stage. As a result, the final product quality could be affected. This section explores how to incorporate those dynamic behaviors into our model.
Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption. To incorporate the condenser and dynamics of vapor flow into the model, consider the mass balance equation for water vapor in a chamber [22]
@@
where $j_{\textrm{w}}$ is the total mass flow rate of water vapor, $j_{\textrm{w,max}}$ is the maximum condenser capacity, $V_{\textrm{c}}$ is the volume of a chamber, and $T_{\textrm{avg}}$ is the average temperature in the chamber assumed to be constant for simplification. The mass flow rate of water vapor can be calculated by
$j_{m,\textrm{w}}=n_{\textrm{vial}}A_{z}N_{w},$ (52)
@@
Here we consider a situation where there is some issue with the condenser, and so the total condenser capacity cannot handle the water vapor produced via sublimation during primary drying, with the parameters given in Table 4. The simulation result shows that there is a pressure increase from the normal operating value of 3 Pa to 20 Pa (Fig. 9A). This pressure increase results from the total vapor flow that exceeds the maximum condenser capacity, leading vapor accumulation in the chamber. When the pressure increases, the driving force for sublimation (also the sublimation flux) decreases. Finally, when the sublimation flux balances with the condenser capacity, the system pressure becomes constant. It is also shown that the drying time increases by about an hour (Fig. 9B), and the product temperature is higher (Fig. 9C).

### v2

@@ Introduction
Thermal radiation is significantly more complicated in terms of modeling. The theories and model equations described below are mainly based on Refs. [31] and [39]; the former discusses general thermal radiation theories, while the latter provides a detailed study on the modeling of thermal radiation in lyophilization. Here we summarize only the key elements needed for modeling thermal radiation in the suspended-vial configuration; more detailed discussion can be found in the aforementioned references. Consider radiation exchange between two diffuse, gray surfaces of finite size denoted as surfaces 1 and 2, respectively. In general, the net radiant energy receiving by surface 1 can be written as
$Q_{\textrm{rad}}=\sigma A_{1}\mathcal{F}\big{(}T_{2}^{4}-T_{1}^{4}\big{)},$ (49)
@@
where $A$ is the surface area, $\mathcal{F}$ is the transfer factor, and the subscripts ‘1’ and ‘2’ denote surfaces 1 and 2, respectively. The transfer factor is dependent on the geometry and material properties (emissivity) of both surfaces. For any two diffuse, gray surfaces that form an enclosure,
$Q_{\textrm{rad}}=\dfrac{\sigma\big{(}T_{2}^{4}-T_{1}^{4}\big{)}}{\dfrac{1-\varepsilon_{1}}{\varepsilon_{1}A_{1}}+\dfrac{1}{A_{1}F_{1-2}}+\dfrac{1-\varepsilon_{2}}{\varepsilon_{2}A_{2}}},$ (50)
@@
where $F_{1-2}$ is the view/shape factor. If surface 1 is surrounded by surface 2 (i.e., $F_{1-2}=1$) and surface 2 is much larger than surface 1, (50) can be simplified as
$Q_{\textrm{rad}}=\varepsilon_{1}\sigma A_{1}\big{(}T^{4}_{2}-T^{4}_{1}\big{)}.$ (51)
@@
In (51), the transfer factor $\mathcal{F}=\varepsilon_{1}$. Next, consider the case where a single vial (surface 1) is surrounded by the chamber walls (surface 2). In this case, (51) becomes
$Q_{\textrm{rad}}=\varepsilon_{\textrm{gl}}\sigma A_{r}\big{(}T^{4}_{\textrm{c}}-T^{4}\big{)},$ (52)
@@
where $\varepsilon_{\textrm{gl}}$ is the emissivity of the glass vial. Here the transfer factor $\mathcal{F}=\varepsilon_{\textrm{gl}}$. The amount of radiative heat for the single-vial case represented by (52) defines the upper bound on the radiative heat that one vial can receive from the chamber walls. This is because, when there are multiple vials, the view factor $F_{1-2}$ is less than 1, and so the radiative heat is shared among the vials. As a result, the transfer factor $\mathcal{F}$ for the multiple-vial case must be less than $\varepsilon_{\textrm{gl}}$. Therefore, (52) needs to be modified to account for interactions between vials.
@@
Typical lyophilization of unit doses always consists of a large number of vials, and so thermal radiation exchange exists not only between vials and chamber walls but also between those multiple vials. In batch lyophilization, vials are arranged as an array consisting of many rows, which results in significant differences in heat transfer conditions between the outer and inner vials. In such cases, the outer vials with higher temperature transfer significant heat to the inner vials while simultaneously exchanging heat with the chamber walls. A rigorous way of modeling this complicated radiation is the radiation network approach, which is comprehensively discussed and demonstrated in the context of lyophilization in Ref. [39]. In the suspended-vial configuration, however, vials are typically aligned in a few rows, e.g., a single row in the equipment proposed by Ref. [7]. As a result, all vials experience nearly the same heat transfer heat condition, resulting in negligible heat transfer between vials. Therefore, instead of using the radiation network approach, we modify (52) to follow (49) as
$Q_{\textrm{rad}}=\sigma A_{r}\mathcal{F}\big{(}T^{4}_{\textrm{c}}-T^{4}\big{)},$ (53)
@@
where $\mathcal{F}$ must be less than $\varepsilon_{\textrm{gl}}$ as described in the previous paragraph. This transfer factor $\mathcal{F}$ is best to be estimated from experimental data but can also be approximated mechanistically, e.g., using (50). In our model, the transfer factor $\mathcal{F}_{\textrm{s3}}$ appears in (6), (32), and (42). If the chamber design and vial configuration are identical for both freezing and drying steps, the same value of $\mathcal{F}_{\textrm{s3}}$ can be used for all equations. A similar analysis can also be done for $\mathcal{F}_{\textrm{s1}}$ in (39) and (46).
@@
When there exists both convection and radiation, it is more convenient to linearize and rewrite the radiation part in the form of Newton’s law of cooling; at the temperature of about 300 K, the error caused by linearization is about 0.1% for the temperature difference of 20 K and 2% for the temperature difference of 100 K (see Section 1.3.2 of Ref. [31]), which is tiny. The linearized equation can then be combined with the convection part, with the corresponding heat transfer coefficient taking into account of both convection and radiation, e.g., (8), (38), and (45). This approximation helps simplify the equations and also facilitates the use of an overall heat transfer coefficient (e.g., (29)) as both convection and radiation parts are written in the form of Newton’s law of cooling.
@@
This section defines the default parameter values for the model developed in Sections 3.2, 3.3, and 3.4. These parameter values are either obtained from the literature or set based on the typical values used in lyophilization. This default set of parameters assumes a complete simulation in which the models for freezing (with controlled nucleation), primary drying, and secondary drying are simulated consecutively. Table LABEL:Tab:DefaultParameters lists the default model parameters; parameter values different from those reported in the table are stated explicitly in each section or case study.
@@
$\rho_{\textrm{f}}$ 937 kg/m3 calculated (see Appendix B)
@@
$T_{\textrm{c}},T_{\textrm{u}}$ 273 (before VISF) K
@@
$R_{\textrm{p0}}$ 1.5$\times$104 m/s
$R_{\textrm{p1}}$ 3.0$\times$107 1/s
$R_{\textrm{p2}}$ 10 1/m
@@
$H$ 7.2$\times$10-3 m calculated (see Appendix B)
$m_{\textrm{s}}$ 1.53$\times$10-4 kg calculated (see Appendix B)
$m_{\textrm{w,0}}$ 2.9$\times$10-3 kg calculated (see Appendix B)
@@
The parameters that should be estimated from experimental data include heat transfer coefficients, mass transfer coefficients and cake resistance, and desorption-related parameters (frequency factor and activation energy). Operating conditions that are not measured, e.g., wall temperatures, could also be estimated from data.
@@
The mechanistic model for continuous lyophilization developed in this article is aimed to be used for various purposes, ranging process design and development to the actual operation (e.g., real-time control, state estimation, and optimization). In continuous operation, the process runs continuously with minimal human intervention, and thus efficient (ideally real-time) simulation and computation are important. Therefore, appropriate numerical methods must be applied. The detailed numerical methods and implementation are given in Appendix C. All simulations, calculations, and results presented in this article were performed and generated using MATLAB R2023a.
The model, data, and MATLAB code used in this work are made available (see the Data Availability section). In addition to the MATLAB implementation, we also provide an equivalent software package in Julia, a high-level, open-source programming language for scientific computing, for the selected simulations.
Results and Discussion
@@
The mechanistic model presented in Section 3 is validated using the experimental data reported in Ref. [7]. Our mdel validation is carried out for each lyophilization step separately, with the specific parameters reported in Table 2.
@@
$T_{\textrm{c}},T_{\textrm{u}}$ 272 K estimated from data in Ref. [7]
@@
$T_{\textrm{c}},T_{\textrm{u}}$ 275 K estimated from data in Ref. [7]
@@
$T_{\textrm{c}},T_{\textrm{u}}$ 285 K estimated from data in Ref. [7]
@@
Figure 5: Model validation for the lyophilization of suspended vials using the experimental data from Ref. [7]. Panel A shows the model prediction and experimental data for the product temperature during the freezing step. Panel B shows the model prediction and experimental data for the product temperature (assumed to be measured at the bottom surface) during the primary drying step. The maximum shelf temperatures for Cases 2a and 2b are 263 and 313 K, respectively. Panel C shows the model prediction and experimental data for the average concentration of bound water during the secondary drying step. The initial concentrations for Cases 3a and 3b are 0.088 and 0.075 kg water/kg solid, respectively.
@@
For the freezing step, the simulated temperature and freezing time agree well with the experimental data (Fig. 5A). The product temperature decreases from its initial value of 280 K to the nucleation temperature of about 263 K, where nucleation starts. Upon the first nucleation, the temperature instantaneously rises up from 263 K to the freezing point of about 272 K. The temperature slowly decreases during the solidification phase for about 30 min, following the freezing-point depression. Subsequently, the temperature decreases faster and becomes equilibrium with the environment (chamber wall and cryogenic gas), with the simulated temperature approaching the equilibrium slightly faster than the actual temperature.
@@
The model accurately predicts the evolution of the product temperature and drying time during primary drying for both low (263 K) and high (313 K) shelf temperatures (Fig. 5B). The temperature rises up rapidly at the beginning and gradually increases at a slower rate toward the end of the process. This is because the driving force for sublimation is initially low, and so most of the heat input is used to increase the temperature. As time progresses, the driving force for sublimation is larger; thus, most of the heat input is used for sublimation. The maximum error in temperature prediction is about 2–3 K, while the error in drying time prediction is about 0.2 h for Case 2a and 0.4 h for Case 2b ($\sim$6%). This error could be partly because of the uncertainty in cake resistance, which results from the stochastic nature of the freezing process.
@@
In secondary drying, the simulated concentration of bound water closely aligns with the experimental data for both concentration levels (0.088 and 0.075 kg water/kg solid). The concentration decreases exponentially from its initial value to the final concentration of 0.01 kg water/kg solid, the threshold set for terminating the simulations. This concentration profile is expected given the linear driving force model is used to describe the desorption process. It is important to note that the data used in Fig. 5C are based on the shelf temperature of 293 K. To properly estimate the values of $f_{\textrm{a}}$ and $E_{\textrm{a}}$, it is recommended to obtain the concentration profiles at various shelf temperatures so that the effects of temperature on the desorption process can be quantified.
Prediction of critical process parameters for a complete continuous lyophilization cycle
@@
One of the important aspects of continuous manufacturing is to ensure that the process is operated smoothly, optimally, and safely without human intervention, to minimize downtime and maximize production. In the context of continuous lyophilization, our model can be used to study the evolution of critical process parameters – e.g., temperature, moisture content – throughout the process at various operating conditions, which can help guide process design and optimization.
@@
This work not only presents the first mechanistic model for continuous lyophilization of suspended vials, but also is one of the very few studies that develops a complete model incorporating all three steps of lyophilization, including freezing, primary drying, and secondary drying. This is important specifically for continuous manufacturing where the entire process should be designed and optimized simultaneously.
@@
With the default model parameters in Table LABEL:Tab:DefaultParameters, a complete continuous lyophilization cycle can be simulated, including (1) freezing with controlled nucleation via VISF, (2) primary drying, and (3) secondary drying (Fig. 6). The time evolution of the product temperature (Fig. 6A), mass of ice (Fig. 6B), and concentration of bound water (Fig. 6C) can be obtained, given the operating pressure (Fig. 6D) and temperature (Fig. 6E). This information could be useful for various purposes. For example, it can guide process and equipment design concerning the chamber sizing and velocity profile of each vial, ensuring that the target residence time is met for each chamber. Simulation results can also be used to help identify pressure and temperature profiles to achieve some specific objectives, e.g., minimization of drying time. It is important to note that the entire simulation shown in Fig. 6 can be computed in less than 1 s on a normal laptop, which makes the model very practical to be implemented for any purposes.
@@
Figure 6: Complete simulation for the suspended-vial continuous lyophilization process showing the time evolution of the (A) product average temperature, (B) total mass of ice, (C) average concentration of bound water, (D) operating pressure, and (E) operating temperature.
Visualizing spatiotemporal data
As discussed in Section 3.1, one of the key considerations during the drying steps is to ensure that the product temperature at any location does not exceed the upper limit. Consequently, spatiotemporal data are needed. Spatial variation of temperature and concentration in the product are dependent on several factors. For instance, temperature gradients are larger when the sample thickness and heat transfer coefficient increase (see the analysis in Appendix A).
[…]
Figure 7: Spatiotemporal evolution of the product temperature and sublimation front during primary drying. Only the frozen region is shown in the figure.
Figure 8: Spatiotemporal evolution of the concentration of bound water during secondary drying. The initial concentration is assumed to vary linearly from 5% (0.05 kg water/solid) at the top to 20% (0.2 kg water/kg solid) at the bottom.
Application to freezing: Optimizing freezing protocols
@@
This section employs the model to provide insights into the freezing step, which usually receives little attention compared to the drying steps. As spontaneous nucleation was already investigated in detail by Refs. [12, 13], controlled nucleation, which is preferable for continuous lyophilization, is considered instead.
The evolution of the product temperature and amount of water during VISF at different pressures predicted by our model is shown in Fig. 9. By varying the pressure profile, it is possible to control the rate of change in the product temperature, hence nucleation. The model can also be used to estimate the amount of water that evaporates during VISF as well as the rate of evaporation. This information could be useful for the development of a freezing protocol for controlled nucleation.
@@
Figure 9: Simulation results showing the (A) evolution of the product during temperature and (B) amount of water during VISF at three different pressures, namely $10^{4}$, 5000, and 1000 Pa. At $t=0$, the pressure decreases linearly from $10^{5}$ Pa to the target pressure in 30 s before being held constant. The nucleation temperature is set to 263 K. External heat transfer with the wall and cryogenic gas is omitted from the model so that only the contribution from VISF is observed. Other model parameters are kept at the default values.
Application to drying: Modifying the model for conventional lyophilization
@@
Although the development of the mechanistic model in this article is for continuous lyophilization, the model can be easily modified for batch/conventional lyophilization. The major modifications are to turn off the thermal radiation part in the model and recalculate/estimate the heat transfer coefficient and cake resistance. In this case study, we consider batch primary drying and secondary drying using the data from Refs. [19] and [4], with the specific parameters reported in Table 3.
@@
$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
@@
$\mathcal{F}_{\textrm{s1}},\mathcal{F}_{\textrm{s3}}$ 0 omit the radiation part
@@
    $T_{\textrm{b}}$ increases linearly at the rate of 0.25 K/min from 228.15 K to 258.15 K (Case 4a) and 268.15 K (Case 4b).
@@
    $T_{\textrm{b}}$ increases linearly at the rate of 0.5 K/min from 264 K to 312 K for Case 5.
Our model provides reliable estimates of the product temperature and sublimation interface position during primary drying at two different shelf temperatures (Fig. 10). The model can also accurately predict the product temperature and concentration of bound water during the secondary drying step (Fig. 11). With proper modifications, our continuous lyophilization model is capable of simulating conventional lyophilization.
Figure 10: Comparison of our model prediction and the experimental/simulation data from Ref. [19] for the primary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and reference simulation data for the sublimation front position. The maximum shelf temperatures for Cases 4a and 4b are 258.15 and 268.15 K, while other conditions are identical.
Figure 11: Comparison of our model prediction and the experimental data from Ref. [4] for the secondary drying step in conventional lyophilization. Panel A shows the model prediction and experimental data for the product temperature at the bottom surface. Panel B shows the model prediction and experimental data for the average concentration of bound water.
@@
Typical process modeling focuses on behaviors of the system during its normal operation where the process is operated steadily under the desired conditions. A better process model should be able to simulate important abnormal conditions, e.g., equipment failure, that could occur due to various unexpected scenarios, which is critical in continuous manufacturing as these abnormal operations could affect the reliability, availability, and maintainability (RAM) as well as the safety of the plant/process.
In lyophilization, one of the crucial design considerations is associated with a condenser. During normal operation, the maximum capacity of a condenser must be higher than the rate of vapor production via sublimation, and choked flow should be avoided [1]. This operational constraint is to ensure that there will not be vapor accumulation in the equipment, which could subsequently lead to in a pressure increase. This scenario may not lead to safety-related issues because the operating pressure in lyophilization is low and the amount of ice/water is generally not high enough to create overpressure. Nevertheless, an increase in the total pressure reduces the driving force for sublimation (see (35)), which then prolongs the primary drying step. As a result, the final product quality could be affected. This section explores how to incorporate those dynamic behaviors into our model.
Most primary drying models, including our model, assume that the system pressure is well controlled, which is a reasonable assumption. To incorporate the condenser dynamics into the model, we consider cases where the total condenser capacity is not enough to condense the water vapor produced via sublimation during primary drying. In such cases, the mass balance for the water vapor in a chamber is [22]
@@
where $j_{\textrm{w}}$ is the total mass flow rate of water vapor resulting from sublimation, $j_{\textrm{w,max}}$ is the maximum condenser capacity, $V_{\textrm{c}}$ is the chamber volume, and $T_{\textrm{avg}}$ is the average temperature in the chamber assumed to be constant for simplification. The mass flow rate of water vapor can be calculated by
$j_{\textrm{w}}=n_{\textrm{vial}}A_{z}N_{w},$ (55)
@@
With the specific parameters given in Table 4, the total pressure increases from the normal operating value of 3 Pa to 20 Pa (Fig. 12A). This pressure increase results from the total vapor flow exceeding the maximum condenser capacity, thus vapor accumulation in the chamber. When the system pressure increases, the driving force for sublimation (also the sublimation flux) decreases. After about 1 h, the system pressure becomes constant, indicating the equalization of the sublimation flux and condenser capacity. In this abnormal operation, the drying time and product temperature are higher than for the normal operation (Figs. 12BC).